NASA is expected to send a 2600-kg satellite 450 km above the earth’s surface.What speed will it have?(Hint: Earth’s mass is 5.97 x 1024 kg)


1
Expert's answer
2021-11-29T11:47:58-0500

Determine the speed taking into consideration that the centripetal force is equal to the gravitational force:


"m\frac{v^2}{R}=G\frac{M_Em}{R^2},\\\space\\
v=\sqrt{\frac{GM}R}=7.6\text{ km/s}."